OpenStudy (anonymous):

Please help quick Solve two-fifths plus one-ninth . Then use estimation to check for reasonableness. Show your work.

OpenStudy (anonymous):

\[\frac{2}{5}+\frac{1}{9}\] \[\frac{2 \times 9}{5 \times 9} + \frac{1 \times 5}{9 \times 5}\]